The perimiter of a rectangle is 150 inches. The length exceeds the width by 29 inches. Find the length and width.
Answer by mananth(16946)

let the with be x
the length = x+29
perimeter = 2x + 2(x+29) = 150
2x+2x+58=150
4x=150-58
4x=92
x= 23 thw width
length = 23 +29 = 52
52+52+23+23= 150